# NeCo[Uncaptioned image]: Improving DINOv2’s spatial representations in 19 GPU hours with Patch Neighbor Consistency

### v1

@@ A.1 Dense Post-Pretraining
Our model is implemented in Python, using Torch [59] and PyTorch Lightning [60].
@@
Our pretraining datasets consist of COCO [56] and ImageNet-100 subset of the original Imagenet [61]. COCO contains approximately 118,000 scene-centric images, whereas ImageNet-100k includes around 100k object-centric images.
@@
Our Data augmentations are the same as in [1]. More specifically, we use: random color-jitter, Gaussian blur, grayscale and multi-crop augmentations. Similarly, the global crop’s resolution is 224x224 and the local crop’s resolution is 96x96, for the all the experiments except when working with Dinov2 where we use 98x98 for local crops. Furthermore, our generated global and local crops have the constraint that they intersect at least by $1\%$ of the original image size.
@@
For our backbone, we employ vision transformers. More specifically, we train on ViT-Small and ViT-Base [46]. Moreover, following [13, 29], we use a student-teacher setup where the teacher weights are updated by the exponential moving average of the student weights.
@@
Following [13], the projection head consists of three linear layers with hidden dimensionality of $2048$, a Gaussian error linear units as activation function [62], and an output dimensionality of $256$.
@@
Following [1], due to the distinction between global and local crops, after projecting the dense spatial output to a lower space, the alignment step is applied on the dense image representations to bring them to a fixed spatial resolution of size of 7x7 during training. This ensures that the local and global crop feature maps have the same size and that they correspond to each other. The alignment is done using region of interest alignment (roi align) [47].
@@
We train both network sizes with a cosine learning rate schedule going down to 00 over $25$ training epochs, except for the ablation studies where we use 10 epochs. The initial projection head learning rate is $1\mathrm{e}{-4}$ for all the experiments, whereas the backbone’s learning rate is $1\mathrm{e}{-5}$, with the exception of being $1\mathrm{e}{-6}$ when applying our method on Dinov2. The exponential moving average for updating the teacher’s weights is adapted with a cosine schedule starting at $0.9995$ and going up to $1$. We use Adam Optimizer [63] with a cosine weight decay schedule.
@@
By default we use the Bitonic Differentiable Sorting Networks [12] and the steepnesses (i.e., inverse temperatures) used for the network are $100$ for the Student and $100$ for the teacher. All the other parameters remain as the default ones; i.e., we use the $logistic_{\phi}$ function with a $\lambda=0.25$ for the interpolation of numbers in the differentiable sorting algorithms.
@@ A.2 Evaluation Setup
The Dense Nearest Neighbor Retrieval Evaluation is a retrieval-based scene understanding evaluation introduced by [10]. Its goal is to assess the scene understanding capabilities of a dense image encoder. It works as follows:
@@
Due to the unavailability of the original implementation by [10], we use the open implementation from [50]. This implementation aligns with the original authors’ description and details, including the use of the ScaNN Library [64] for efficient nearest neighbor retrieval. We adhere to the setup from the Hummingbird Model authors [10] for our experiments. We use a memory size of 10,240,000 and configure ScaNN with 30 nearest neighbors, consistent with the evaluation of the Hummingbird model on this memory size.
The final results are reported as mean Intersection over Union (mIoU) on four different fractions of two datasets: Pascal VOC 2012 [65] and ADE20K [54]. The sub-sampling factors are 1, 8, 64, or 128. For factors greater than 1, results are averaged over five different seeds. These dataset subsets are created by uniformly and randomly selecting a unique set of images from the training split, ensuring an approximately equal number of distinct images for each annotation label. For example, for the 1/128 fraction of the Pascal VOC 2012 dataset, we would collect around 83 images, ensuring each of the 20 labels (excluding the background) appears in at least 4 different images in the subset.
Overclustering
For the Overclustering experiment, following [1], we run $K$-Means (using faiss [66]) on all spatial tokens from our backbone (i.e., with the projection head discarded) for a given dataset. We then group the clusters to the ground-truth classes of the dataset by applying greedy matching to the pixel-level precision and then run Hungarian matching [49] on the combined cluster maps, which makes the evaluation metric permutation-invariant [67]. We use a crop size of 448x448 for the input images, and overclustering is applied on downsampled 100x100 masks in order to speed up the Hungarian matching. The final results are reported as an average of mean Intersection over Union (mIoU) over five different seeds on four different datasets: COCO-Thing and COCO-Stuff [56], Pascal VOC 2012 [65], and ADE20K [54].
@@
For linear segmentation, we closely follow the setup from Leopart [1]. Concisely, we take 448x448 images, encode them with our backbone to get the spatial outputs, apply bilinear interpolation to match the mask resolution, and finally apply a linear head to obtain the segmentation predictions. These predictions are then compared with the ground truth segmentation masks and trained via cross-entropy loss.
@@
Moreover, we train and evaluate linear heads on four versions of datasets: Pascal VOC 2012 [65], subsets of COCO-Thing and COCO-Stuff (explained in Appendix A), and ADE20K [54].
@@
Following the evaluation setup from [9], we finetune our backbones and the transformer-based decoder from Segmenter [48] in an end-to-end manner. We use the Segmenter implementation available within the MMSegmentation Library [68].
The performance metric used here is the mIoU score, reported on four different datasets: Pascal Context [52], Pascal VOC 2012 [65], COCO-Stuff 164K [56], and ADE20K [54]. The crop size used is 512×512. For the DINOv2 model and our method on it, we apply zero padding around the image of 512×512 to bring it to the size of 518×518.
The remaining configurations follow [9]. For the ADE20K and COCO-Stuff 164K datasets, we use 160k iterations, and for Pascal VOC 2012 and Pascal Context, we use 80k iterations, all with an eta_min of $0.1\cdot lr$. We use the Adam optimizer [63] and for each pretraining method and dataset, we experiment with four different learning rates ($8\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, $8\times 10^{-6}$) before reporting the highest mIoU score.
@@
To better evaluate the scene understanding abilities of our method, we also evaluate it using the Fully Unsupervised Semantic Segmentation Evaluation method [1]. This evaluation consists of two parts: Cluster-based Foreground Extraction (CBFE) and Overclustering with Community Detection (CD).
The CBFE clusters the spatial outputs of a model over a dataset and assigns each cluster as background (bg) or foreground (fg). The separation of foreground and background clusters is facilitated by attention maps from a Vision Transformer, which provide cues for the fg/bg distinction. We construct the final hard fg-bg assignment by averaging the attention heads, applying Gaussian filtering with a 7x7 kernel size, and retaining $70\%$ of the attention mass to obtain the final binary mask. The rest of the configurations remain the same as the original setup [1].
The CD metric [1] exploits local co-occurrence statistics among clusters to identify and categorize objects. This approach uses no labels for categorizing semantic parts; it simply finds local co-occurrence of clusters in an image by utilizing an information-theoretic definition of network communities. Our configurations for the CD evaluation remain the same as in Leopart [1].
We use the implementation from Leopart [1] and apply CBFE and CD on the non-augmented (train) split of Pascal VOC 2012 [65], and evaluate on its full validation set. For CD, we report the best results over 10 seeds obtained from a hyper-parameter search, leading to our best parameters for CD+CBFE: weight_threshold = 0.07, markov_time = 1.2, and k_community = 189.
@@ B.2 Full Clustering Tables
Table 7: Clustering evaluation performance. K-means with various clustering granularity $K$ is applied to the spatial features obtained from different feature extractors on two datasets. The resulting cluster maps are matched to the ground truth by Hungarian matching [49], and the intersection is reported in mIoU.
@@
DINO [13] ViT-S/16 21M 4.3 13.9 17.3 5.4 18.8 19.2
iBOT [14] ViT-S/16 21M 4.4 23.8 31.1 7.6 26.6 28.0
CrOC [4] ViT-S/16 21M 3.4 16.4 20.0 4.9 14.7 18.1
TimeT [2] ViT-S/16 21M 12.2 43.6 46.2 17.5 42.7 44.6
DINOv2R [15] ViT-S/14 21M 12.2 46.7 49.5 12.3 38.9 41.2
CrIBo [9] ViT-S/16 21M 18.3 51.3 54.5 14.5 46.0 48.3
@@ B.3 Full Visual In-Context Learning Tables
We show the results shown by Figure 2 in Table 8. To provide a more comprehensive comparison, we also evaluate our method against SelfPatch [69], a finetuning approach with a similar aim of enhancing dense representations.
Table 8: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, ADE20k and Pascal VOC. The retrieved cluster maps are compared with the ground truth using Hungarian matching [49], and their mIoU score is reported.
@@
DINO ViT-S/16 21M 9.5 11.0 15.0 17.9 26.4 30.5 41.3 47.9
SelfPatch ViT-S/16 21M 10.0 10.9 14.7 17.7 28.4 32.6 43.2 50.1
CrOC ViT-S/16 21M 8.7 10.8 15.2 17.3 34.0 41.8 53.8 59.5
TimeT ViT-S/16 21M 12.1 14.1 18.9 23.2 38.1 43.8 55.2 61.3
Leopart ViT-S/16 21M 12.9 14.8 19.6 23.9 44.6 49.7 58.4 63.4
CrlBo ViT-S/16 21M 14.6 17.3 22.7 26.6 53.9 59.9 66.9 71.2
DINOv2R ViT-S/14 21M 19.6 22.8 30.1 35.9 53.0 57.9 68.2 73.8
NeCo ViT-S/14 21M 23.7 27.2 34.0 39.8 66.5 70.3 76.3 78.9
MAE ViT-B/16 85M 10.0 11.3 15.4 18.6 3.5 4.1 5.6 6.9
DINO ViT-B/16 85M 11.5 13.5 18.2 21.5 33.1 37.7 49.8 56.5
Leopart ViT-B/16 85M 14.6 16.8 21.8 26.7 50.1 54.7 63.1 68.2
@@
CrlBo ViT-B/16 85M 15.9 18.4 24.4 28.4 55.9 61.8 69.2 73.1
DINOv2R ViT-B/14 85M 22.1 25.8 33.2 38.7 51.8 58.9 70.6 76.2
NeCo ViT-B/14 85M 29.1 33.7 39.2 44.2 67.0 71.4 77.5 82.2
B.4 Extra Ablations

### v2

@@ A.1 Dense Post-Pretraining
Our model is implemented in Python, using Torch (Paszke et al., 2019) and PyTorch Lightning (Falcon & team, 2019).
@@
Our pretraining datasets consist of COCO (Caesar et al., 2018) and ImageNet-100 subset of the original Imagenet (Russakovsky et al., 2015). COCO contains approximately 118,000 scene-centric images, whereas ImageNet-100k includes around 100k object-centric images.
@@
Our Data augmentations are the same as in Ziegler & Asano (2022). More specifically, we use: random color-jitter, Gaussian blur, grayscale and multi-crop augmentations. Similarly, the global crop’s resolution is 224x224 and the local crop’s resolution is 96x96, for the all the experiments except when working with Dinov2 where we use 98x98 for local crops. Furthermore, our generated global and local crops have the constraint that they intersect at least by $1\%$ of the original image size.
@@
For our backbone, we employ vision transformers. More specifically, we train on ViT-Small and ViT-Base (Dosovitskiy et al., 2020). Moreover, following Caron et al. (2021); Grill et al. (2020), we use a student-teacher setup where the teacher weights are updated by the exponential moving average of the student weights.
@@
Following Caron et al. (2021), the projection head consists of three linear layers with hidden dimensionality of $2048$, a Gaussian error linear units as activation function (Hendrycks & Gimpel, 2016), and an output dimensionality of $256$.
@@
Following Ziegler & Asano (2022), due to the distinction between global and local crops, after projecting the dense spatial output to a lower space, the alignment step is applied on the dense image representations to bring them to a fixed spatial resolution of size of 7x7 during training. This ensures that the local and global crop feature maps have the same size and that they correspond to each other. The alignment is done using region of interest alignment (roi align) (He et al., 2017).
@@
We train both network sizes with a cosine learning rate schedule going down to 00 over $25$ training epochs, except for the ablation studies where we use 10 epochs. The initial projection head learning rate is $1\mathrm{e}{-4}$ for all the experiments, whereas the backbone’s learning rate is $1\mathrm{e}{-5}$, with the exception of being $1\mathrm{e}{-6}$ when applying our method on Dinov2. The exponential moving average for updating the teacher’s weights is adapted with a cosine schedule starting at $0.9995$ and going up to $1$. We use Adam Optimizer (Kingma & Ba, 2017) with a cosine weight decay schedule.
@@
By default we use the Bitonic Differentiable Sorting Networks (Petersen et al., 2021) and the steepnesses (i.e., inverse temperatures) used for the network are $100$ for the Student and $100$ for the teacher. All the other parameters remain as the default ones; i.e., we use the $logistic_{\phi}$ function with a $\lambda=0.25$ for the interpolation of numbers in the differentiable sorting algorithms.
@@ A.2 Evaluation Setup
The Dense Nearest Neighbor Retrieval Evaluation is a retrieval-based scene understanding evaluation introduced by Balazevic et al. (2023). Its goal is to assess the scene understanding capabilities of a dense image encoder. It works as follows:
@@
Due to the unavailability of the original implementation by Balazevic et al. (2023), we use the open implementation from Pariza et al. (2024). This implementation aligns with the original authors’ description and details, including the use of the ScaNN Library (Guo et al., 2020) for efficient nearest neighbor retrieval. We adhere to the setup from the Hummingbird Model authors (Balazevic et al., 2023) for our experiments. We use a memory size of 10,240,000 and configure ScaNN with 30 nearest neighbors, consistent with the evaluation of the Hummingbird model on this memory size.
The final results are reported as mean Intersection over Union (mIoU) on four different fractions of two datasets: Pascal VOC 2012 (Everingham et al., ) and ADE20K (Zhou et al., 2017). The sub-sampling factors are 1, 8, 64, or 128. For factors greater than 1, results are averaged over five different seeds. These dataset subsets are created by uniformly and randomly selecting a unique set of images from the training split, ensuring an approximately equal number of distinct images for each annotation label. For example, for the 1/128 fraction of the Pascal VOC 2012 dataset, we would collect around 83 images, ensuring each of the 20 labels (excluding the background) appears in at least 4 different images in the subset.
Overclustering.
For the Overclustering experiment, following Ziegler & Asano (2022), we run $K$-Means (using faiss (Johnson et al., 2019)) on all spatial tokens from our backbone (i.e., with the projection head discarded) for a given dataset. We then group the clusters to the ground-truth classes of the dataset by applying greedy matching to the pixel-level precision and then run Hungarian matching (Kuhn, 1955) on the combined cluster maps, which makes the evaluation metric permutation-invariant (Ji et al., 2019). We use a crop size of 448x448 for the input images, and overclustering is applied on downsampled 100x100 masks in order to speed up the Hungarian matching. The final results are reported as an average of mean Intersection over Union (mIoU) over five different seeds on four different datasets: COCO-Thing and COCO-Stuff (Caesar et al., 2018), Pascal VOC 2012 (Everingham et al., ), and ADE20K (Zhou et al., 2017).
@@
For linear segmentation, we closely follow the setup from Leopart (Ziegler & Asano, 2022). Concisely, we take 448x448 images, encode them with our backbone to get the spatial outputs, apply bilinear interpolation to match the mask resolution, and finally apply a linear head to obtain the segmentation predictions. These predictions are then compared with the ground truth segmentation masks and trained via cross-entropy loss.
@@
Moreover, we train and evaluate linear heads on four versions of datasets: Pascal VOC 2012 (Everingham et al., ), subsets of COCO-Thing and COCO-Stuff (explained in Appendix A), and ADE20K (Zhou et al., 2017).
@@
Following the evaluation setup from Lebailly et al. (2023), we finetune our backbones and the transformer-based decoder from Segmenter (Strudel et al., 2021) in an end-to-end manner. We use the Segmenter implementation available within the MMSegmentation Library (MMSegmentation Contributors, 2020).
The performance metric used here is the mIoU score, reported on four different datasets: Pascal Context (Everingham et al., 2010), Pascal VOC 2012 (Everingham et al., ), COCO-Stuff 164K (Caesar et al., 2018), and ADE20K (Zhou et al., 2017). The crop size used is 512×512. For the DINOv2 model and our method on it, we apply zero padding around the image of 512×512 to bring it to the size of 518×518.
The remaining configurations follow Lebailly et al. (2023). For the ADE20K and COCO-Stuff 164K datasets, we use 160k iterations, and for Pascal VOC 2012 and Pascal Context, we use 80k iterations, all with an eta_min of $0.1\cdot lr$. We use the Adam optimizer (Kingma & Ba, 2017) and for each pretraining method and dataset, we experiment with four different learning rates ($8\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, $8\times 10^{-6}$) before reporting the highest mIoU score.
@@
To better evaluate the scene understanding abilities of our method, we also evaluate it using the Fully Unsupervised Semantic Segmentation Evaluation method (Ziegler & Asano, 2022). This evaluation consists of two parts: Cluster-based Foreground Extraction (CBFE) and Overclustering with Community Detection (CD).
The CBFE clusters the spatial outputs of a model over a dataset and assigns each cluster as background (bg) or foreground (fg). The separation of foreground and background clusters is facilitated by attention maps from a Vision Transformer, which provide cues for the fg/bg distinction. We construct the final hard fg-bg assignment by averaging the attention heads, applying Gaussian filtering with a 7x7 kernel size, and retaining $70\%$ of the attention mass to obtain the final binary mask. The rest of the configurations remain the same as the original setup (Ziegler & Asano, 2022).
The CD metric (Ziegler & Asano, 2022) exploits local co-occurrence statistics among clusters to identify and categorize objects. This approach uses no labels for categorizing semantic parts; it simply finds local co-occurrence of clusters in an image by utilizing an information-theoretic definition of network communities. Our configurations for the CD evaluation remain the same as in Leopart (Ziegler & Asano, 2022).
We use the implementation from Leopart (Ziegler & Asano, 2022) and apply CBFE and CD on the non-augmented (train) split of Pascal VOC 2012 (Everingham et al., ), and evaluate on its full validation set. For CD, we report the best results over 10 seeds obtained from a hyper-parameter search, leading to our best parameters for CD+CBFE: weight_threshold = 0.07, markov_time = 1.2, and k_community = 189.
@@ B.2 Full Clustering Tables
Table 8: Clustering evaluation performance. K-means with various clustering granularity $K$ is applied to the spatial features obtained from different feature extractors on two datasets. The resulting cluster maps are matched to the ground truth by Hungarian matching (Kuhn, 1955), and the intersection is reported in mIoU.
@@
DINO (Caron et al., 2021) ViT-S/16 21M 4.3 13.9 17.3 5.4 18.8 19.2
iBOT (Zhou et al., 2022) ViT-S/16 21M 4.4 23.8 31.1 7.6 26.6 28.0
CrOC (Stegmüller et al., 2023) ViT-S/16 21M 3.4 16.4 20.0 4.9 14.7 18.1
TimeT (Salehi et al., 2023) ViT-S/16 21M 12.2 43.6 46.2 17.5 42.7 44.6
DINOv2R (Oquab et al., 2023) ViT-S/14 21M 12.2 46.7 49.5 12.3 38.9 41.2
CrIBo (Lebailly et al., 2023) ViT-S/16 21M 18.3 51.3 54.5 14.5 46.0 48.3
@@ B.3 Full Visual In-Context Learning Tables
We show the results shown by Figure 2 in Table 9. To provide a more comprehensive comparison, we also evaluate our method against SelfPatch (Yun et al., 2022), a finetuning approach with a similar aim of enhancing dense representations.
Table 9: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, ADE20k and Pascal VOC. The retrieved cluster maps are compared with the ground truth using Hungarian matching (Kuhn, 1955), and their mIoU score is reported.
@@
DINO ViT-S/16 21M 9.5 11.0 15.0 17.9 26.4 30.5 41.3 48.7
SelfPatch ViT-S/16 21M 10.0 10.9 14.7 17.7 28.4 32.6 43.2 50.8
CrOC ViT-S/16 21M 8.7 10.8 15.2 17.3 34.0 41.8 53.8 60.5
TimeT ViT-S/16 21M 12.1 14.1 18.9 23.2 38.1 43.8 55.2 62.3
Leopart ViT-S/16 21M 12.9 14.8 19.6 23.9 44.6 49.7 58.4 64.5
CrlBo ViT-S/16 21M 14.6 17.3 22.7 26.6 53.9 59.9 66.9 72.4
DINOv2R ViT-S/14 21M 19.6 22.8 30.1 35.9 53.0 57.9 68.2 75.0
NeCo ViT-S/14 21M 23.7 27.2 34.0 39.8 66.5 70.3 76.3 80.2
MAE ViT-B/16 85M 10.0 11.3 15.4 18.6 3.5 4.1 5.6 7.0
DINO ViT-B/16 85M 11.5 13.5 18.2 21.5 33.1 37.7 49.8 57.3
Leopart ViT-B/16 85M 14.6 16.8 21.8 26.7 50.1 54.7 63.1 69.5
@@
CrlBo ViT-B/16 85M 15.9 18.4 24.4 28.4 55.9 61.8 69.2 74.2
DINOv2R ViT-B/14 85M 22.1 25.8 33.2 38.7 51.8 58.9 70.6 77.3
NeCo ViT-B/14 85M 29.1 33.7 39.2 44.2 67.0 71.4 77.5 83.5
B.4 Object detection with ViTDet
We report the performance of ViTDet (Li et al., 2022) on COCO dataset for ViT-S with DINO and DINOv2R backbones and compare them with the NeCo  finetuned versions. Our method improves DINO and DINOv2R backbones, specially DINO, increasing the validation Box AP slightly from 42.9 to 43.0 and the Mask AP from 38.6 to 38.7.
Table 10: Object detection performance with ViTDet. Although NeCo  is trained only for 19 GPU hours, it can still improve the performance of DINO backbones on average across the specified measures.
Backbone Epochs Val Box AP Val Mask AP
Dino 12 42.9 38.6
+ NeCo 12 43.0 38.7
Dinov2 12 42.5 36.7
+ NeCo 12 42.6 36.7
B.5 Applying NeCo  to vision-language foundation models
We present the results for CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) on linear segmentation and visual in-context learning in Table 12 and Table 11. As shown in the tables above, NeCo  improves the performance of both CLIP and SigLIP by approximately 12% to 37% across various benchmarks. These results demonstrate that NeCo  is not limited to vision foundation models but can also be effectively applied to vision-language models.
Table 11: In-context scene understanding benchmark (mIoU). Dense nearest neighbor retrieval performance is reported across various training data proportions on ADE20k and Pascal VOC datasets.
ADE20K Pascal VOC
Method Backbone Params 1/128 1/64 1/8 1/1 1/128 1/64 1/8 1/1
CLIP ViT-B/16 85M 5.8 6.5 8.7 11.3 25.3 27.8 33.4 33.9
+ NeCo ViT-B/16 85M 17.7 19.8 22.9 24.2 62.8 63.5 65.1 66.2
SIGLIP ViT-B/16 85M 6.0 7.1 9.0 10.6 25.3 27.8 32.2 33.9
+ NeCo ViT-B/16 85M 15.9 18.4 20.7 21.9 60.9 62.0 62.5 63.1
Table 12: Linear segmentation performance. A linear segmentation head is trained on top of the frozen spatial features obtained from different feature extractors. We report the mIoU scores achieved on the validation sets of 5 different datasets.
Method Backbone Params Pascal VOC ADE20K COCO-Stuff COCO-Things Cityscapes
CLIP ViT-B/16 85M 44.3 13.8 43.1 42.0 27.7
+NeCo ViT-B/16 85M 68.2 25.8 56.1 48.9 41.0
SigLIP ViT-B/16 85M 44.6 15.9 36.2 46.0 32.2
+NeCo ViT-B/16 85M 70.1 29.4 55.3 69.9 42.0
B.6 Generalizing NeCo  to broader tasks and datasets
we report the performance of various backbones, including DINOv2R finetuned with the NeCo loss, on out-of-distribution datasets such as Cityscapes (Cordts et al., 2016)(semantic segmentation) and NYUd (Couprie et al., 2013)(monocular depth estimation). As demonstrated in 13(a) and 13(c), NeCo  consistently enhances the generalization of all backbones on the Cityscapes dataset. Furthermore, even for tasks that diverge from semantic segmentation, such as depth estimation, NeCo  reduces the DINOv2R error by around 2%. These findings highlight that NeCo not only maintains but improves the generalization capability of DINOv2R features across diverse tasks. For completeness, we also include results on (Markus Gerke, 2014), showing that NeCo consistently enhances the performance, even on out-of-distribution datasets.
Table 13: Performance comparison on the Cityscapes, NYUd, and Vaihingen datasets.
(a) Linear segmentation on Cityscapes
Pretraining Original + NeCo
DINO-S/16 40.7 43.7
iBot-S/16 43.0 44.8
TimeT-S/16 42.2 42.8
Leopard-S/16 44.5 45.5
Dinov2R-S/14 48.9 52.9
(b) Linear segmentation on Vaihingen
Pretraining Original + NeCo
CLIP-B/16 25.7 28.8
SigLIP-B/16 26.9 28.3
Dinov2R-B/14 34.2 35.9
(c) Linear depth prediction on NYUd
Backbone Pretraining RMSE
ViT-S Dinov2 0.460
Dinov2R 0.456
+ NeCo 0.453
ViT-B Dinov2 0.412
Dinov2R 0.410
+ NeCo 0.397
B.7 Extra Ablations